When two full moons appear in the same month,the second full moon is called a blue moon. On average, 10 blue moons occur every 25 years. Find the number of blue moons that are likely to occur in the next 41 years.

Answer by stanbon(75887)

When two full moons appear in the same month,the second full moon is called a blue moon. On average, 10 blue moons occur every 25 years. Find the number of blue moons that are likely to occur in the next 41 years.
--------------------
x/41 = 10/25
x = 41(10/25 = 41(2/5) = 82/5 = 16 2/5
Round down to get x = 16 blue moons in the next 41 years
==================================
Cheers,
Stan H.